how many milliliters of 3.5M NaCl must be diluted to prepare 250ml of 0.45 M NaCl
1
Expert's answer
2015-12-08T10:15:46-0500
N(NaCl) = 0.25 * 0.45 = 0.1125 (mol);
V(NaCl) = 0.1125 / 3.5 = 0.032 (l);

Answer: 32 ml